OpenStudy (.whitedragon.):

A cylinder has a radius of 12 m and a height of 9 m. What is the exact volume of the cylinder?

OpenStudy (whpalmer4):

\[v=\pi r^2 h\] Plug in the numbers, leave \(\pi\) as a symbol

OpenStudy (.whitedragon.):

1296? pie squared

OpenStudy (anonymous):

4069.44

OpenStudy (whpalmer4):

@Mr.Man420 That's not an exact value, because \(\pi\) has an unending set of digits, so multiplying it with an integer (such as 12*12*9) must also give an unending set of digits.

OpenStudy (whpalmer4):

\(\pi\) is the Greek letter "pi", not "pie" :-) How are you getting \(\pi^2\) in your answer?

OpenStudy (anonymous):

lol

OpenStudy (whpalmer4):

\[V = \pi r^2h = \pi (12\text{ m})^2(9\text{ m}) = \pi * 12*12*9 \text{ m}^3 =\]

OpenStudy (anonymous):

so how could you get an exact answer if its unending?

OpenStudy (whpalmer4):

By leaving \(\pi\) as a symbol. \(3\pi\) is an exact quantity. You just can't write down a string of digits that expresses it exactly.

OpenStudy (.whitedragon.):

closest estimation is 3.14

OpenStudy (whpalmer4):

No, that's not the closest estimation. That's about the least closest estimation :-)

OpenStudy (.whitedragon.):

:p

OpenStudy (whpalmer4):

Here's a closer estimation, for example: \(\pi \approx \)3.14159265358979323846264338327950288419716939937510582097494459230781\ 6406286208998628034825342117067982148086513282306647093844609550582231\ 7253594081284811174502841027019385211055596446229489549303819644288109\ 7566593344612847564823378678316527120190914564856692346034861045432664\ 8213393607260249141273724587006606315588174881520920962829254091715364\ 3678925903600113305305488204665213841469519415116094330572703657595919\ 5309218611738193261179310511854807446237996274956735188575272489122793\ 8183011949129833673362440656643086021394946395224737190702179860943702\ 7705392171762931767523846748184676694051320005681271452635608277857713\ 4275778960917363717872146844090122495343014654958537105079227968925892\ 3542019956112129021960864034418159813629774771309960518707211349999998\ 3729780499510597317328160963185950244594553469083026425223082533446850\ 3526193118817101000313783875288658753320838142061717766914730359825349\ 0428755468731159562863882353787593751957781857780532171226806613001927\ 876611195909216420199

OpenStudy (.whitedragon.):

What is the exact volume of the cylinder?

OpenStudy (anonymous):

Math, so confuseing...

OpenStudy (whpalmer4):

@.WhiteDragon. haven't seen your correct answer for the original problem yet...

OpenStudy (.whitedragon.):

1296m pi squared

OpenStudy (whpalmer4):

You said that already. That isn't correct.

OpenStudy (.whitedragon.):

What is the exact volume of the cylinder?

OpenStudy (whpalmer4):

\[V = \pi r^2h = \pi (12\text{ m})^2(9\text{ m}) = \pi * 12*12*9 \text{ m}^3 =\]

OpenStudy (.whitedragon.):

well then what is it

OpenStudy (.whitedragon.):

1296 pi m squared

OpenStudy (whpalmer4):

No. Why do you keep saying that? How do you multiply \[\large\pi*12*12*9 \text{ m}^3 \]and get the answer you keep supplying?

OpenStudy (whpalmer4):

Is \(\pi\) being squared anywhere?

OpenStudy (whpalmer4):

Units of volume have dimensions \[\text{[length]}^3\]so the units are incorrect as well

OpenStudy (.whitedragon.):

well then help

OpenStudy (whpalmer4):

All you have to do is carefully and correctly multiply the quantities I put in front of you. You keep saying "squared" — what is making you think that you should say that?

OpenStudy (whpalmer4):

Here's a hint: If you have a square with sides \(3\text{ m}\), the area is given by \[A = s^2 = (3\text{ m})(3\text{ m}) = 3*3*\text{m}*\text{m} = 9\text{ m}^2\] If you have a cube with sides of the same length, the volume is given by \[V = s^3 = (3\text{ m})(3\text{ m})(3\text{ m}) = 3*3*3*\text{m}*\text{m}*\text{m} = 27\text{ m}^3\] The units are not plucked out of the air — they are a direct result and part of the calculation. If you have the wrong units, you may well have made additional mistakes elsewhere in the calculation. That's one of the reasons why understanding how to use and manipulate units is important.

OpenStudy (.whitedragon.):